You roll a fair six-sided die twice. Find the probability of rolling a 1 the first time and a number greater than 5 the second time.
Answer by mananth(16949)

Probability of rolling a 1is 1/6, then probabulity of rolling a 5 is 1/6.
1/6 * 1/6 = 1/36
These are independent events
The probability of first rolling a 5, then a 3, is 1/6 * 1/6 = 1/36